OpenStudy (fatimab2000):

Will Give Medal !!!!!! Sabrina is using an online calculator to calculate the outputs f(n) for different inputs n. The ordered pairs below show Sabrina's inputs and the corresponding outputs displayed by the calculator. (1, 6), (2, 11), (3, 16), (4, 21) Which of the following functions best represents the rule that the calculator uses to display the outputs? a. f(n) = 5n - 1 b. f(n) = 5n + 1 c. f(n) = 4n + 3 d. f(n) = 4n - 3

OpenStudy (anonymous):

it is the second one

OpenStudy (fatimab2000):

THx

OpenStudy (anonymous):

do you have other questions

OpenStudy (fatimab2000):

The function f(x) is shown below: f(x) = (4)^(x + 5) The function f(x) is shifted to the right by 13 units. Which of the following best represents the new function? f(x) = (4)^(x - 8) f(x) = (4)^(x + 13) f(x) = (4)^(x + 18) f(x) = (4)^(x - 13)

OpenStudy (anonymous):

Either A or C because choice A corresponds to shifting the function down by 13 units while choice C corresponds to shifting it up by 13 units. Check the question again?

OpenStudy (fatimab2000):

ok thx

OpenStudy (anonymous):

that's all

OpenStudy (fatimab2000):

yup

OpenStudy (anonymous):

ok bye nice helping you